OpenStudy (anonymous):

can you help me get the y' of this equation? x^1/2 + y^1/2 = a^1/2

OpenStudy (anonymous):

By y' do you mean dy/dx?

OpenStudy (anonymous):

yep. dy/dx

OpenStudy (anonymous):

or inverse

OpenStudy (anonymous):

I would use implicit differentiation. I get y'=(1/2)(x^(-1/2)+1/(a^(1/2)-x^(1/2))) although it may be wrong.

OpenStudy (anonymous):

Wait that's wrong, hold on...

OpenStudy (anonymous):

the book says the answer is \[(-y ^1/2) / (x^1/2)\]

OpenStudy (anonymous):

but i don't know how the answer arrived to be like that.

OpenStudy (anonymous):

\[\sqrt{x}+\sqrt{y}=\sqrt{a}\] \[\frac{1}{2\sqrt{x}}+\frac{1}{2\sqrt{y}}y^'=0\] \[\frac{1}{2\sqrt{y}}y^'=-\frac{1}{2\sqrt{x}}\] \[y'=-\frac{\sqrt{y}}{\sqrt{x}}\]

OpenStudy (anonymous):

Yep that's right. I got a minus sign wrong and then resubstituted to get rid of the y. First diff implicitly to get 1/(2x^(1/2))+(y')/(2^(1/2))=0 etc. What the person above said, lol

OpenStudy (anonymous):

why did the variable a turned to be equal to zero?

OpenStudy (anonymous):

in this case a is assumed to be a constant, whose derivative is 0

OpenStudy (anonymous):

because a^(1/2) is independent of x

OpenStudy (anonymous):

i see. thanks for the help. :) i'll give you two a medal.